a 0.750 m solution of ascorbic acid has a pH of 2.11. calculate the ka of ascorbic acid
1
Expert's answer
2020-06-02T14:08:06-0400

"pH = (pK_a -logc)/2 \implies pK_a=2pH+logc"

"pK_a= 4.22-0.125=4.095=-logK_a"

"\implies K_a=10^{-4.095}= 8.035*10^{-5}"